OpenStudy (anonymous):

The expression 4n – 1 is a rule to describe a certain sequence. What are the first three terms of the sequence? A 3, 7, 11 B 4, 8, 12 C 3, 6, 9 D 4, 7, 11

hartnn (hartnn):

to get the 1st term , put n=1 in 4n-1. what u get ?

OpenStudy (anonymous):

D?

hartnn (hartnn):

lol! no... put 1 in 4n-1 means 4(1)-1 = 4-1 = ?

OpenStudy (anonymous):

I got 7 :/

hartnn (hartnn):

4-1 = 7 ?? O.o

hartnn (hartnn):

i thought 4-1 = 3

hartnn (hartnn):

and 8-1=7 and 12-1=11